STAT 230 Exam 2 with Solution, Spring 2011
American University of Beirut · Introduction to Probability and Random Variables · 2 pages · Dr. Mohammad Kobeissi · with solution
More STAT 230 files
- Exam 1 · Summer 2016ExamView
- Exam 2 Solution · Fall 2016Exam · with solutionView
- Exam 2 Solution · Fall 2016Exam · with solutionView
- Exam 1 · Summer 2016ExamView
- Midterm with Answers · Summer 2015Exam · with solutionView
- Midterm Summer 2015 · Summer 2015ExamView
- Exam 1 · Summer 2013ExamView
- Exam 1 With Solution · Summer 2013Exam · with solutionView
Show the text of this exam
American University of Beirut
STAT 230
Introduction to Probability and Random Variables
Spring 2010
quiz # 2 solution
1. It is known that 30% of all calls coming into
a telephone exchange are long-distance calls.
Find the probability that the first long-distance
call is among the first 50 coming calls.
Let A ={event the first long-distance call in
among the first 50}; then A={no long-distance
call is among the first 50}, and
P(A) = 1 − P(A) = 1 − (0.7)50 = 0.99
2. A certain typing agency employs two typists.
The average number of errors per article is 3
when typed by the first typist and 4.5 when
typed by the second typist. If your article is
equally likely to be typed by either typist, find
the probability that it will have exactly one
error.
The average number of error by typist 1 is X ⇝
P(3), and the average number of error by typist
2 is Y ⇝ P(4.5).
Let A ={event article have exactly one error}
P(A) = 1
2P(X = 1) + 1
2P(Y = 1) = 1
2(0.149 +
0.0499) = 0.099 (from the Poisson table)
3. The probability that a machine produces a de-
fective item is 0.01. Each item is checked as it
is produced. Approximate the probability that
at least 3 items are defective among the next
120 items checked.
Let X be the number of defective items among
the 120; X ⇝ b(120, 0.01).
X can be approximated by a Poisson distribu-
tion Y with parameter λ = 1.2
P(X ≥ 3) ≃ P(Y ≥ 3) = 0.12 (from the Pois-
son table)
4. The number of years a radio functions is ex-
ponentially distributed with parameter λ = 1
8.
If Sam buys a used radio, find the probability
that it will be working after an additional 8
years.
Let X be the lifetime (in years) of a radio; X ⇝
E(1/8), and
f(x) = 1
8 e−x/8
0 < x < ∞
Let y be the lifetime of the radio when Sam
bought it,
P(X > 8 + y|X > y) = P(X > y + 8)
P(X > y)
=
� ∞
y+8
1
8 e−x/8 dx
� ∞
y
1
8 e−x/8 dx = e−(y+8)/8
e−y/8
= e−1
(you may also say it directly P(X > 8) after
mentioning that the exponential distribution is
memory less)
5. A town has 2 fire engines operating indepen-
dently. The probability that a specific engine
is available when needed is 0.96. Find the prob-
ability that at least one fire engine is available
when needed.
Let A ={event at least one fire engine is avail-
able}
P(A) = 1 − P(A) = 1 − (0.04)2 = 0.9984
6. You arrive at a bus stop at 10 o’clock, know-
ing that bus will arrive at some time uniformly
distributed between 10 and 10:30. If at 10:15
the bus has not arrived yet, what is the prob-
ability that you will have to wait at least an
additional 10 minutes ?
The arriving time X of the bus can be modeled
by a uniform distribution on the interval (0, 1)
(here we consider that 10 o’clock is 0 and 10:30
is 1)
P(X > 5/6|X > 1/2) = P(X>5/6)
P(X>1/2) = 1/6
1/2 = 1/3
7. A roulette wheel has 18 red cases and 20 black
cases. Suppose that you continue to make $5
bet on red until you win 4 of these bets.
i Find the probability that you place a total
of 9 bets.
Let X be the number of bets until you win
4 bets; X has a negative binomial distribu-
tion with parameters r = 4 and p = 18/38;
then
P(X = 9) = C3
8( 20
38)5( 18
38)4
ii Find your expected winning when you stop.
The expected number of bets until you win
4 is r/p = 76/9, hence you are expected to
play 76/9 games; out of theses games 4 are
wining games and the rest is losing.
Your expected wining is then given by
4 × 5 − ( 76
9 − 4) × 5 = −20
9 $
8. Let X be a random variable with pdf
f(x) = 3x2 ,
0 < x < 1
i Find F(x), the cdf of X.
F(x) =
0
x < 0
x3
0 < x < 1
1
x > 1
ii Find E(Xn).
E(Xn) =
� 1
0
3xn+2 dx =
3
n + 3
9. Sam and Jad play a game in which Sam’s chance
of winning is 2/3. Suppose that if Jad win a
game then Sam must pay him 1 $, and that if
Sam win a game then Jad must pay him 0.75 $.
Find the expected gain of Jad in a series of 50
such games, supposedly independent.
Let X be the number of games won by Jad
among the 50; X ⇝ b(50, 1
3).
The expected number of games that Sam will
win is E(X) = 50/3, hence the expected wining
of Sam is
1 × 50
3 − 0.75 × 100
3 = −25
3